Assuming the angular resolution for the Hubble Telescope to be 2.86 ✕ 10−7 rad, what is the smallest detail that could be observed on Venus when it is closest to Earth? (Give the size of the smallest detail in meters. Assume the distance from the Hubble Telescope to the surface of Venus is approximately 3.82 ✕ 1010 m.)

(Answer in m.)
Answer #1

Given,

The angular resolution, = 2.86 * 10-7 rad

The distance from telescope to the surface of Venus, r = 3.82 * 1010 m

The smallest detail that is observed on Venus when closest to Earth, s = r *

s = 3.82 * 1010 * 2.86 * 10-7

= 10.925 * 103 m